### v1

@@ 3 Score-Based MMSE Denoising
$\mathbb{E}\left[\mathbf{x}|\tilde{\mathbf{x}}\right]$ $=\tilde{\mathbf{x}}+\sigma^{2}\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}),$ (11)
$\mathrm{Cov}\left[\mathbf{x}|\tilde{\mathbf{x}}\right]$ $=\sigma^{2}\mathbf{I}+\sigma^{4}\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}).$ (12)
Theorem 1 combines Tweedie’s formula [26, 11] for retrieving the posterior mean and its second-order generalization for retrieving the posterior variance [22, 12]. We provide a unified proof in Appendix A. Remarkably, Theorem 1 establishes the theoretical foundation for using score functions to achieve empirical Bayes-optimal denoising of AWGN observations. Once the score functions are known, we can apply Theorem 1 to efficiently implement the MMSE denoiser in (9) and (10). We discuss how to obtain the first- and second-order score estimates, i.e., $\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})$ and $\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}})$, via the denoising score matching technique [31, 29] in the following.
4 First- and Second-Order Score Matching
4.1 First-Order Score Matching
For a given $\sigma$, we aim to learn a first-order score model $\mathbf{s}_{\bm{\theta}}(\cdot,\sigma):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ parameterized by $\bm{\theta}$ with the following score matching objective:
$\min_{\bm{\theta}}~{}\mathbb{E}_{p(\tilde{\mathbf{x}})}\left[\left\|\mathbf{s}_{\bm{\theta}}(\tilde{\mathbf{x}},\sigma)-\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})\right\|^{2}_{2}\right].$ (13)
The above formulation is shown equivalent to the denoising score matching objective [31, 29], given by
$\min_{\bm{\theta}}~{}\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\left[\left\|\mathbf{s}_{\bm{\theta}}(\tilde{\mathbf{x}},\sigma)-\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right\|^{2}_{2}\right].$ (14)
@@ 4.1 First-Order Score Matching
$\min_{\bm{\theta}}~{}\ell_{1}(\bm{\theta};\sigma)\coloneq\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\left[\left\|\mathbf{s}_{\bm{\theta}}(\tilde{\mathbf{x}},\sigma)+\frac{\tilde{\mathbf{x}}-\mathbf{x}}{\sigma^{2}}\right\|^{2}_{2}\right].$ (15)
@@
$\min_{\bm{\theta}}~{}\mathcal{L}_{1}\big{(}\bm{\theta};\{\sigma_{i}\}_{i=1}^{L}\big{)}\coloneq\frac{1}{L}\sum_{i=1}^{L}\lambda_{1}(\sigma_{i})\ell_{1}(\bm{\theta};\sigma_{i}),$ (16)
@@
4.2 Second-Order Score Matching
For a given $\sigma$, we aim to learn a second-order score model $\mathbf{S}_{\bm{\phi}}(\cdot,\sigma):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N\times N}$ parameterized by $\bm{\phi}$ with the following score matching objective:
$\min_{\bm{\phi}}~{}\mathbb{E}_{p(\tilde{\mathbf{x}})}\left[\left\|\mathbf{S}_{\bm{\phi}}(\tilde{\mathbf{x}},\sigma)-\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}})\right\|^{2}_{F}\right].$ (17)
@@ 4.2 Second-Order Score Matching
$\min_{\bm{\phi}}~{}$ $\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\bigg{[}\Big{\|}\mathbf{S}_{\bm{\phi}}(\tilde{\mathbf{x}},\sigma)-\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)^{\sf T}+\frac{\mathbf{I}}{\sigma^{2}}\Big{\|}_{F}^{2}\bigg{]},$ (18)
where $\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\coloneq\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})+\frac{\tilde{\mathbf{x}}-\mathbf{x}}{\sigma^{2}}$. Please refer to Appendix B for the detailed derivation. The objective in (18) relies on the ground-truth of the first-order score function, which is in general not accessible. In practice, we replace $\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})$ by the learned $\mathbf{s}_{\bm{\theta}}(\tilde{\mathbf{x}},\sigma)$ for efficient computation. The objective in (18) then becomes
$\min_{\bm{\phi}}~{}$ $\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\bigg{[}\Big{\|}\mathbf{S}_{\bm{\phi}}(\tilde{\mathbf{x}},\sigma)-\hat{\mathbf{b}}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\hat{\mathbf{b}}(\mathbf{x},\tilde{\mathbf{x}},\sigma)^{\sf T}+\frac{\mathbf{I}}{\sigma^{2}}\Big{\|}_{F}^{2}\bigg{]},$ (19)
where $\hat{\mathbf{b}}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\coloneq\mathbf{s}_{\bm{\theta}}(\tilde{\mathbf{x}},\sigma)+\frac{\tilde{\mathbf{x}}-\mathbf{x}}{\sigma^{2}}$. It is proved in [19] that with this replacement, the second-order score model has the error-bounded property if the first-order score matching error is bounded.
As in Line 12 of Algorithm 1, only the trace of the posterior covariance matrix is needed, rather than the covariance matrix itself. By noting in (12) the simple relation between the posterior covariance matrix and the second-order score function, it suffices to only match the trace of the second-order score function for posterior variance evaluation. The simplified objective is expressed as
$\min_{\bm{\phi}}~{}\ell_{2}(\bm{\phi};\sigma)$ :-$\coloneq\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\bigg{[}\bigg{|}\mathrm{tr}\big{(}\mathbf{S}_{\bm{\phi}}(\tilde{\mathbf{x}},\sigma)\big{)}-\left\|\hat{\mathbf{b}}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\right\|_{2}^{2}+\frac{N}{\sigma^{2}}\bigg{|}^{2}\bigg{]}.:$ (20)
We combine (20) for all $\sigma\in\{\sigma_{i}\}_{i=1}^{L}$ to obtain the following unified objective for training the second-order score model:
$\min_{\bm{\phi}}~{}\mathcal{L}_{2}\big{(}\bm{\phi};\{\sigma_{i}\}_{i=1}^{L}\big{)}\coloneq\frac{1}{L}\sum_{i=1}^{L}\lambda_{2}(\sigma_{i})\ell_{2}(\bm{\phi};\sigma_{i}),$ (21)
where $\lambda_{2}(\sigma_{i})$ is the weighting factor depending on $\sigma_{i}$. In practice, one can first train a first-order score model, and then freeze it when applied to second-order score learning. Alternatively, the two score models can be trained both at once by disabling gradient backpropagation from the second-order training objective (21) to the first-order score model.
5 STMP Algorithm
Algorithm 2 STMP Algorithm
@@ 5 STMP Algorithm
11:     $\mathbf{x}_{\sf B}^{\textit{post}}=\mathbf{x}_{\sf B}^{\textit{pri}}+v_{\sf B}^{\textit{pri}}\mathbf{s}_{\bm{\theta}}\big{(}\mathbf{x}_{\sf B}^{\textit{pri}},\sqrt{v_{\sf B}^{\textit{pri}}}\big{)}$
12:     $v_{\sf B}^{\textit{post}}=\frac{(v_{\sf B}^{\textit{pri}})^{2}}{N}\mathrm{tr}\Big{(}\mathbf{S}_{\bm{\phi}}\big{(}\mathbf{x}_{\sf B}^{\textit{pri}},\sqrt{v_{\sf B}^{\textit{pri}}}\big{)}\Big{)}+v_{\sf B}^{\textit{pri}}$
@@
$\mathbf{x}_{\sf B}^{\textit{post}}$ $=\mathbf{x}_{\sf B}^{\textit{pri}}+v_{\sf B}^{\textit{pri}}\mathbf{s}_{\bm{\theta}}\big{(}\mathbf{x}_{\sf B}^{\textit{pri}},\sqrt{v_{\sf B}^{\textit{pri}}}\big{)},$ (22)
$v_{\sf B}^{\textit{post}}$ $=\frac{(v_{\sf B}^{\textit{pri}})^{2}}{N}\mathrm{tr}\Big{(}\mathbf{S}_{\bm{\phi}}\big{(}\mathbf{x}_{\sf B}^{\textit{pri}},\sqrt{v_{\sf B}^{\textit{pri}}}\big{)}\Big{)}+v_{\sf B}^{\textit{pri}}.$ (23)
As shown in Lines 5 and 6 of Algorithm 2, the LMMSE estimator requires the computation of a different matrix inverse at each iteration with prohibitive complexity. This can be avoided by the pre-calculation of the singular value decomposition (SVD) of $\mathbf{A}$ [25]. Moreover, in compressive imaging, the rows of $\mathbf{A}$ are typically randomly selected from an orthogonal matrix (e.g., DCT matrix). This results in a diagonal matrix to be inverted at each iteration, which significantly reduces the computational cost.
6 State Evolution Analysis
The asymptotic MSE performance of the STMP algorithm can be accurately predicted by a set of SE equations. Due to the limited representation capability of a neural network, the score-based denoiser may not be perfectly matched with the MMSE performance at each iteration. The LMMSE estimator can also be substituted with other linear estimators, such as matched filter and pseudo-inverse [20]. Therefore, we adopt a more general notation to represent the linear estimator and the denoiser as
$\mathbf{x}_{\sf A}^{\textit{post}}=\mathbf{g}_{\sf A}(\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}}),~{}~{}~{}~{}\mathbf{x}_{\sf B}^{\textit{post}}=\mathbf{g}_{\sf B}(\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}}).$ (26)
We make the following assumption on the posterior variances.
Assumption 1.
Assume that the posterior variances perfectly match the linear estimator and the denoiser under the practical setting in (26), i.e.,
$v_{\sf A}^{\textit{post}}$ $=\frac{1}{N}\mathbb{E}\left[(\mathbf{g}_{\sf A}(\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}})-\mathbf{x})^{\sf T}(\mathbf{g}_{\sf A}(\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}})-\mathbf{x})\right],$ (27)
$v_{\sf B}^{\textit{post}}$ $=\frac{1}{N}\mathbb{E}\left[(\mathbf{g}_{\sf B}(\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}})-\mathbf{x})^{\sf T}(\mathbf{g}_{\sf B}(\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}})-\mathbf{x})\right].$ (28)
Similar to [13], define the MSE function of the linear estimator under AWGN measurement as
$\mathrm{mse}_{\sf A}(v_{\sf A}^{\textit{pri}},\tau_{\sf A})\coloneq\lim_{N\rightarrow\infty}\frac{1}{N}\mathbb{E}\left[\left\|\mathbf{g}_{\sf A}(\mathbf{r}_{\sf A},v_{\sf A}^{\textit{pri}})-\mathbf{x}\right\|^{2}_{2}\right],$
$\mathbf{r}_{\sf A}=\mathbf{x}+\mathcal{N}(\mathbf{0},\tau_{\sf A}\mathbf{I}),~{}~{}~{}~{}\mathbf{y}=\mathbf{A}\mathbf{x}+\mathcal{N}(\mathbf{0},\delta_{0}^{2}\mathbf{I}),$ (29)
and define that of the denoiser as
$\mathrm{mse}_{\sf B}(v_{\sf B}^{\textit{pri}},\tau_{\sf B})\coloneq\lim_{N\rightarrow\infty}\frac{1}{N}\mathbb{E}\left[\left\|\mathbf{g}_{\sf B}(\mathbf{r}_{\sf B},v_{\sf B}^{\textit{pri}})-\mathbf{x}\right\|^{2}_{2}\right],$
$\mathbf{r}_{\sf B}=\mathbf{x}+\mathcal{N}(\mathbf{0},\tau_{\sf B}\mathbf{I}).$ (30)
Furthermore, we define the sensitivity functions of the linear estimator and the denoiser respectively as
$\mathrm{sf}_{\sf A}(v_{\sf A}^{\textit{pri}},\tau_{\sf A})$ :-$\coloneq\lim_{N\rightarrow\infty}\frac{1}{N}\mathrm{tr}\left(\frac{\partial\mathbf{g}_{\sf A}(\mathbf{r}_{\sf A},v_{\sf A}^{\textit{pri}})}{\partial\mathbf{r}_{\sf A}}\right),:$ (31)
$\mathrm{sf}_{\sf B}(v_{\sf B}^{\textit{pri}},\tau_{\sf B})$ :-$\coloneq\lim_{N\rightarrow\infty}\frac{1}{N}\mathrm{tr}\left(\frac{\partial\mathbf{g}_{\sf B}(\mathbf{r}_{\sf B},v_{\sf B}^{\textit{pri}})}{\partial\mathbf{r}_{\sf B}}\right).:$ (32)
Theorem 2.
Under Assumptions 1, 1, 2, and 3, the SE of the STMP algorithm at the $k$-th iteration is established by the following scalar equations:
$v_{{\sf A},k}^{\textit{post}}$ $=v_{{\sf A},k}^{\textit{pri}}\mathrm{sf}_{\sf A}(v_{{\sf A},k}^{\textit{pri}},\tau_{{\sf A},k}),$ (33)
$v_{{\sf B},k}^{\textit{pri}}$ $=v_{{\sf A},k}^{\textit{ext}}=\frac{v_{{\sf A},k}^{\textit{pri}}v_{{\sf A},k}^{\textit{post}}}{v_{{\sf A},k}^{\textit{pri}}-v_{{\sf A},k}^{\textit{post}}},$ (34)
$\tau_{{\sf B},k}$ $=\frac{(v_{{\sf A},k}^{\textit{pri}})^{2}\mathrm{mse}_{\sf A}(v_{{\sf A},k}^{\textit{pri}},\tau_{{\sf A},k})-(v_{{\sf A},k}^{\textit{post}})^{2}\tau_{{\sf A},k}}{(v_{{\sf A},k}^{\textit{pri}}-v_{{\sf A},k}^{\textit{post}})^{2}},$ (35)
$v_{{\sf B},k}^{\textit{post}}$ $=v_{{\sf B},k}^{\textit{pri}}\mathrm{sf}_{\sf B}(v_{{\sf B},k}^{\textit{pri}},\tau_{{\sf B},k}),$ (36)
$v_{{\sf A},k+1}^{\textit{pri}}$ $=v_{{\sf B},k}^{\textit{ext}}=\frac{v_{{\sf B},k}^{\textit{pri}}v_{{\sf B},k}^{\textit{post}}}{v_{{\sf B},k}^{\textit{pri}}-v_{{\sf B},k}^{\textit{post}}},$ (37)
$\tau_{{\sf A},k+1}$ $=\frac{(v_{{\sf B},k}^{\textit{pri}})^{2}\mathrm{mse}_{\sf B}(v_{{\sf B},k}^{\textit{pri}},\tau_{{\sf B},k})-(v_{{\sf B},k}^{\textit{post}})^{2}\tau_{{\sf B},k}}{(v_{{\sf B},k}^{\textit{pri}}-v_{{\sf B},k}^{\textit{post}})^{2}},$ (38)
in which $\tau_{{\sf A},k}$ and $\tau_{{\sf B},k}$ denote the true noise variances in the inputs of module A and module B, respectively.
Theorem 2 restates the main result presented in [13]. For the relevant assumptions, please refer to Appendix C. We initialize the equations with $\tau_{{\sf A},0}=v_{{\sf A},0}^{\textit{pri}}=\mathbb{E}\big{[}\big{\|}\mathbf{x}_{\sf A}^{\textit{pri}}-\mathbf{x}\big{\|}_{2}^{2}\big{]}$. At the $k$-th iteration, the asymptotic MSE of STMP is predicted by $\mathrm{mse}_{\sf B}(v_{{\sf B},k}^{\textit{pri}},\tau_{{\sf B},k})$.
The SE equations can be further simplified when the score-based networks act as an MMSE denoiser. The simplified SE equations coincide with the fixed-point characterization of the Bayes-optimal MSE predicted by the replica method [25]. This implies that STMP can potentially achieve the Bayes-optimal MSE performance with a perfectly learned MMSE denoiser.
7 Experiments
In this section, we evaluate the performance of the proposed method based on the FFHQ $256\times 256$ dataset [16].
Training and inference configurations.

### v2

@@ III Score-Based MMSE Denoising
[…]
Theorem 1 establishes the theoretical foundation for using score functions to achieve empirical Bayes-optimal denoising of AWGN observations. Once the score functions are known, we can apply Theorem 1 to efficiently implement the MMSE denoiser in (10) and (11). We discuss how to obtain the first- and second-order score functions via the denoising score matching technique [19, 13] in the following.
IV First- and Second-Order Score Matching
IV-A First-Order Score Matching
For a given $\sigma$, we aim to learn a first-order score model $\mathbf{s}_{\boldsymbol{\theta}}(\cdot,\sigma):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ parameterized by $\boldsymbol{\theta}$ with the following score matching objective:
$\min_{\boldsymbol{\theta}}~\mathbb{E}_{p(\tilde{\mathbf{x}})}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\tilde{\mathbf{x}},\sigma)-\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})\right\|^{2}_{2}\right].$ (16)
The above formulation is shown equivalent to the denoising score matching objective [19], given by
$\min_{\boldsymbol{\theta}}~\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\tilde{\mathbf{x}},\sigma)-\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right\|^{2}_{2}\right].$ (17)
@@ IV-A First-Order Score Matching
$\min_{\boldsymbol{\theta}}~\ell_{1}(\boldsymbol{\theta};\sigma)\triangleq\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\tilde{\mathbf{x}},\sigma)+\frac{\tilde{\mathbf{x}}-\mathbf{x}}{\sigma^{2}}\right\|^{2}_{2}\right].$ (18)
@@
$\min_{\boldsymbol{\theta}}~\mathcal{L}_{1}\big(\boldsymbol{\theta};\{\sigma_{i}\}_{i=1}^{L}\big)\triangleq\frac{1}{L}\sum_{i=1}^{L}\lambda_{1}(\sigma_{i})\ell_{1}(\boldsymbol{\theta};\sigma_{i}),$ (19)
@@
IV-B Second-Order Score Matching
For a given $\sigma$, we aim to learn a second-order score model $\mathbf{S}_{\boldsymbol{\phi}}(\cdot,\sigma):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N\times N}$ parameterized by $\boldsymbol{\phi}$ with the following score matching objective:
$\min_{\boldsymbol{\phi}}~\mathbb{E}_{p(\tilde{\mathbf{x}})}\left[\left\|\mathbf{S}_{\boldsymbol{\phi}}(\tilde{\mathbf{x}},\sigma)-\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}})\right\|^{2}_{F}\right].$ (20)
@@ IV-B Second-Order Score Matching
$\min_{\boldsymbol{\phi}}$~ $\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\bigg[\Big\|\mathbf{S}_{\boldsymbol{\phi}}(\tilde{\mathbf{x}},\sigma)$
$~~~~~~~~~~~~~~~-\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)^{\sf T}+\frac{\mathbf{I}}{\sigma^{2}}\Big\|_{F}^{2}\bigg],$ (21)
[…]
As indicated in (11), only the trace of the posterior covariance matrix is needed, rather than the covariance matrix itself. By noting in (15) the simple relation between the posterior covariance matrix and the second-order score function, it suffices to only match the trace of the second-order score function for posterior variance evaluation. The simplified objective is expressed as
$\min_{\boldsymbol{\phi}}~\ell_{2}(\boldsymbol{\phi};\sigma)$ $\triangleq\mathbb{E}_{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}\bigg[\bigg|\mathrm{tr}\big(\mathbf{S}_{\boldsymbol{\phi}}(\tilde{\mathbf{x}},\sigma)\big)$
$~~~~~~~~~~~~~~~-\left\|\hat{\mathbf{b}}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\right\|_{2}^{2}+\frac{N}{\sigma^{2}}\bigg|^{2}\bigg],$ (22)
where $\hat{\mathbf{b}}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\triangleq\mathbf{s}_{\boldsymbol{\theta}}(\tilde{\mathbf{x}},\sigma)+(\tilde{\mathbf{x}}-\mathbf{x})/\sigma^{2}$. We combine (IV-B) for all $\sigma\in\{\sigma_{i}\}_{i=1}^{L}$ to obtain the following unified objective for training the second-order score model:
$\min_{\boldsymbol{\phi}}~\mathcal{L}_{2}\big(\boldsymbol{\phi};\{\sigma_{i}\}_{i=1}^{L}\big)\triangleq\frac{1}{L}\sum_{i=1}^{L}\lambda_{2}(\sigma_{i})\ell_{2}(\boldsymbol{\phi};\sigma_{i}),$ (23)
where $\lambda_{2}(\sigma_{i})$ is the weighting factor depending on $\sigma_{i}$.
In practice, one can first train a first-order score model, and then freeze it when applied to second-order score learning. Alternatively, the two score models can be trained both at once by disabling gradient backpropagation from the second-order training objective (23) to the first-order score model.
V STMP Algorithm
$\mathbf{x}_{\sf B}^{\textit{post}}$ $=\mathbf{x}_{\sf B}^{\textit{pri}}+v_{\sf B}^{\textit{pri}}\mathbf{s}_{\boldsymbol{\theta}}\big(\mathbf{x}_{\sf B}^{\textit{pri}},\sqrt{v_{\sf B}^{\textit{pri}}}\big),$ (24)
$v_{\sf B}^{\textit{post}}$ $=\frac{(v_{\sf B}^{\textit{pri}})^{2}}{N}\mathrm{tr}\Big(\mathbf{S}_{\boldsymbol{\phi}}\big(\mathbf{x}_{\sf B}^{\textit{pri}},\sqrt{v_{\sf B}^{\textit{pri}}}\big)\Big)+v_{\sf B}^{\textit{pri}}.$ (25)
As shown in Lines 5 and 6 of Algorithm 1, the LMMSE estimator requires the computation of a different matrix inverse at each iteration with prohibitive complexity. This can be avoided by the pre-calculation of the singular value decomposition (SVD) of $\mathbf{A}$ [16]. Moreover, in compressive imaging, the rows of $\mathbf{A}$ are typically randomly selected from an orthogonal matrix (e.g., discrete cosine transform (DCT) matrix). This results in a diagonal matrix to be inverted at each iteration, which significantly reduces the computational cost.
@@ V STMP Algorithm
11:  $\mathbf{x}_{\sf B}^{\textit{post}}=\mathbf{x}_{\sf B}^{\textit{pri}}+v_{\sf B}^{\textit{pri}}\mathbf{s}_{\boldsymbol{\theta}}\big(\mathbf{x}_{\sf B}^{\textit{pri}},\sqrt{v_{\sf B}^{\textit{pri}}}\big)$
12:  $v_{\sf B}^{\textit{post}}=\frac{(v_{\sf B}^{\textit{pri}})^{2}}{N}\mathrm{tr}\Big(\mathbf{S}_{\boldsymbol{\phi}}\big(\mathbf{x}_{\sf B}^{\textit{pri}},\sqrt{v_{\sf B}^{\textit{pri}}}\big)\Big)+v_{\sf B}^{\textit{pri}}$
@@
Algorithm 1 STMP Algorithm
VI Experiments